Question 04 08 Marks
Price (Rupees per Ice-cream cup) Quantity demanded (Ice-cream cup per month) Quantity Supplied (Ice-cream cup per month)
140 500 1500
120 750 1200
100 1000 1000
80 1250 750
60 1500 600
40 1750 300

a) Plot the demand and supply curves and show the equilibrium price and quantity. Also, briefly explain if law of demand and supply holds true in this case. (2 Marks)
b) Is there surplus or shortage in the market at price Rs.40? At price Rs.120? (2 Marks)
c) Suppose if the price of frozen yogurt decreases, elucidate the change it might cause on the ice-cream market given the consumption behavior and also illustrates the same through using 2-dimensional space. In your opinion what type of elasticity best describes this situation? (3 marks)
d) Suppose the price of milk decreases from the equilibrium price. Explain what will happen to the supply curve and also indicate the same on the graph. (1mark)
1
Expert's answer
2020-11-30T16:29:19-0500
"Solution"

a) The equilibrium price "=100" and equilibrium quantity "=1000". The law of demand with all factors held constant has a negative relationship between price and quantity demanded. When price falls from $140 to $40, quantity demanded will rise from 500 to 1,750.

Yes, the law of supply also holds, since there is a positive relationship between price and quantity supplied. When price fall from $140 to $40, quantity supplied falls from 1500 to 300



b) When

"p=\$40, There\ is\ a\ shortage\ as\ Q_d>Q_s\\
Shortage=\ Q_d-Q_s=1,750-300=1,450\ cups\\
When\ p=\$120,\ there\ is\ surplus\ as\ Q_d<Q_s\\
Surplus=Q_s-Q_d=1,200-750=450\ cups"

c) Frozen yogurt and ice cream are substitute goods. Therefore, a decrease in price of frozen yogurt will lead to a decrease in demand for ice cream. Due to this, the demand curve will shift to the left, which is a cross price elasticity of demand.

This is depicted in the graph below.



d) Decrease in price of milk will lead to an increase in supply of milk. Since milk is an input for ice cream, the supply curve will shift to the right as shown below.